If f(x)=x^2+5, find f(a+h)-f(a). Thank you for any help you can provide.
Answer by stanbon(75887)

If f(x)=x^2+5, find f(a+h)-f(a).
f(a+h)=(a+h)^2+5=a^2+2ah+h^2+5
f(a)= a^2+5
f(a+h)-f(a)= 2ah+h^2
Cheers,
Stan H.